Please help me solve this algebra 2 question:
The Absolute value of 5-4i

Answer by venugopalramana(3286)

ABSOLUTE VALUE OF A COMPLEX NUMBER Z = A+iB IS GIVEN BY
|Z|=SQUARE ROOT OF (A^2+B^2)...HENCE FOR Z = 5-4i.WE GET
|Z|=SQUARE ROOT OF (5^2+(-4)^2)= SQUARE ROOT OF (25+16)= SQUARE ROOT OF (41)